### v1

# Impact of general relativistic accretion on primordial black holes

Abstract
We demonstrate that general relativistic corrections to the accretion of relativistic matter onto primordial black holes (PBHs) can significantly enhance their mass growth during the early Universe. Contrary to previous Newtonian treatments, our analysis reveals that PBH masses can increase by an order of magnitude before evaporation, leading to substantial modifications of their lifetime and cosmological imprints. We quantify the resulting shifts in the minimum PBH mass constrained by Big Bang Nucleosynthesis (BBN), the revised lower bound for PBHs surviving today, and the dark matter parameter space allowed by PBH evaporation. Furthermore, we show that the enhanced accretion alters the high-frequency gravitational wave spectrum from PBH evaporation, potentially within the reach of future detectors. Our results provide a comprehensive, relativistically consistent framework to delineate the role of PBHs in early-universe cosmology and dark matter phenomenology.
Primordial black holes, Reheating, Superradiance, Dark matter, Hawking radiation.
I Introduction
Black holes (BHs) have been at the forefront of theoretical physics, bridging the realms of classical and quantum gravity. Traditionally, their phenomenological studies revolved around the gravitational capture of surrounding matter through the process of accretion in astrophysical environments Hoyle & Lyttleton (1939); 1952MNRAS.112..195B . However, a quantum mechanical perspective reveals that BHs can also emit particles through Hawking radiation Hawking:1975vcx , an insight that reshaped the understanding of the very foundational nature of gravity and paved the way into its cosmological implications.
The groundbreaking discovery of gravitational waves by the LIGO and Virgo collaborations in 2016 LIGOScientific:2016aoc further propelled BHs into the spotlight, intensifying their relevance in observational cosmology. In this realm, primordial black holes (PBHs) stand out as a potential key to unlocking several cosmological puzzles. These PBHs, formed from the gravitational collapse of overdensities in the early universe, span a wide mass range — from sub-gram scales to solar masses Carr & Hawking (1974); Carr (1975); Carr:2020gox ; Auffinger:2022khh ; Khlopov:2008qy . Depending on their mass, PBHs can undergo both accretion and evaporation during cosmic evolution, influencing various phenomena such as dark matter (DM) production Masina:2020xhk ; Hooper:2019gtx ; Baker:2022rkn ; Cheek:2021odj ; Calabrese:2021src ; Chaudhuri:2023aiv , baryogenesis Baumann:2007yr ; Fujita:2014hha ; Datta:2020bht ; Hook:2014mla ; Barman:2022gjo , and cosmic microwave background (CMB) distortions Tashiro:2008sf ; Acharya:2020jbv ; Ricotti:2007au ; Ali-Haimoud:2016mbv ; Piga:2022ysp ; Poulin:2017bwe ; Serpico:2020ehh ; Ziparo:2022fnc to name a few. Accretion is one of the pivotal processes of PBHs evolution in the early universe Michel (1972); Harada:2004pf ; Harada:2004pe ; Lora-Clavijo:2013aya ; Mack:2006gz ; DeLuca:2020fpg ; DeLuca:2021pls ; DeLuca:2023bcr ; Yang:2022puh ; Yuan:2023bvh ; Zhang:2023rnp . While early studies on PBH accretion primarily focused on Bondi-Hoyle Ricotti:2007jk ; Tabasi:2021cxo ; Tabasi:2022fap ; Stojkovic:2004hz ; Jangra:2024sif ; Nayak:2009wk ; Park:2010yh ; Wang:2023qxj ; Manzoni:2023gon ; khlopov:hal-03129467 models in the Newtonian framework, the general relativistic effect could be particularly important for PBHs that are embedded in a hot and dense environment of relativistic plasma in the early universe.
Motivated by this, we present a reformulation of the PBH accretion process within a general relativistic framework—a regime that, to the best of our knowledge, remains largely unexplored, and investigate its impact on the evolution of PBHs for the first time. Our analysis shows that relativistic corrections to the Bondi accretion model significantly enhance the accretion rate, leading to substantial mass growth of ultralight PBHs prior to their evaporation via Hawking radiation. By numerically solving the coupled equations governing accretion and evaporation in an expanding Universe, we derive updated constraints on the initial PBH mass spectrum, refine bounds on their abundance, and assess their viability as dark matter candidates. In addition, we predict a measurable shift in the high-frequency gravitational wave spectrum due to accretion-induced mass evolution. This unified treatment of PBH dynamics incorporating both evaporation and relativistic accretion provides a more accurate and testable framework for PBH based dark matter scenarios in light of current cosmological data.
II General relativistic accretion
We present the basic formalism of general relativistic accretion in an axisymmetric, asymptotically flat BH spacetime, following the approach outlined in Dihingia:2018tlr , where spacetime is given by,
$ds^{2}=g_{tt}dt^{2}+g_{rr}dr^{2}+g_{\theta\theta}d\theta^{2}+g_{\phi\phi}d\phi^{2}+2g_{t\phi}d\phi dt.$ (1)
The horizon is defined as $g^{rr}=1/g_{rr}=0$ and the metric coefficients are functions of $(r,\theta)$. In this background, the dynamics of an ideal fluid is described by the conservation equations, $\nabla_{\nu}T^{\mu\nu}=0~{};~{}\nabla_{\mu}j^{\mu}=0$, where the indices $\mu$ and $\nu$ run from 00 to $3$. The symbol $\nabla_{\mu}$ denotes the covariant derivative. Recall that any ideal fluid of energy density $\rho$, pressure $p$, and mass density $\tilde{\rho}$ is characterized by the energy momentum tensor $T^{\mu\nu}$, and the particle number current $j^{\mu}$, which are expressed as,
$T^{\mu\nu}=(\rho+p)u^{\mu}u^{\nu}+pg^{\mu\nu}~{}~{};~{}~{}j^{\mu}=\tilde{\rho}u^{\mu}.$ (2)
Here, $u^{\mu}$ is the four velocity in the lab frame supplemented by the condition $u_{\mu}u^{\mu}=-1$. Defining a projection operator as $h^{i}_{\mu}=\delta^{i}_{\mu}+u^{i}u_{\mu}$, that satisfies $h^{i}_{\mu}u^{\mu}=0$, we project the energy momentum conservation equation into a three vector equation as,
$h^{i}_{\mu}(\nabla_{\nu}T^{\mu\nu})=(\rho+p)u^{\nu}(\nabla_{\nu}u^{i})+(g^{i\nu}+u^{i}u^{\nu})(\partial_{\nu}p)=0,$ (3)
and a scalar equation as,
$u_{\mu}(\nabla_{\nu}T^{\mu\nu})=u^{\mu}\left(h\partial_{\mu}\tilde{\rho}-\partial_{\mu}\rho\right)=0.$ (4)
In equation (3), the roman index assumes $i=1,2,3$. Conventionally, the evaluation of the accreting matter is described in terms of the quantity defined in co-rotating frame of the fluid considering the following series of variable transformations: Azimuthal velocity is defined as $v_{\phi}^{2}=(u^{\phi}u_{\phi})/(-u^{t}u_{t})$ with the associated Lorentz factor $\gamma_{\phi}=1/\sqrt{(1-v^{2}_{\phi})}$; the polar velocity is defined as $v_{\theta}^{2}=\gamma^{2}_{\phi}(u^{\theta}u_{\theta})/(-u^{t}u_{t})$ with the associated Lorentz factor $\gamma_{\theta}=1/\sqrt{(1-v^{2}_{\theta})}$, and finally the radial velocity is defined as $v^{2}=\gamma^{2}_{\phi}\gamma^{2}_{\theta}(u^{r}u_{r})/(-u^{t}u_{t})$ with the associated radial Lorentz factor $\gamma_{v}=1/\sqrt{(1-v^{2})}$. All detailed equations formulated using the co-rotating fluid variables are provided in the reference Dihingia:2018tlr . However, for our present study we focus on spherical accretion and assume $v_{\theta}=v_{\phi}=0$. Expanding the Eq. (3) for $i=r$, we get,
$v\gamma^{2}_{v}\frac{dv}{dr}+\frac{1}{h\tilde{\rho}}\frac{dp}{dr}+\frac{d\Phi^{\text{eff}}}{dr}=0.$ (5)
Here, $\Phi^{\text{eff}}~{}[=\frac{1}{2}\ln(-g_{tt})]$ is the general relativistic effective potential and $h~{}[=\rho+p)/\tilde{\rho}]$ denotes the enthalpy of the flow. In this work, we consider Schwarzschild BH where $g_{tt}=-g^{rr}=-f(r),~{}g_{t\phi}=0,~{}g_{\theta\theta}=g_{\phi\phi}/\sin^{2}\theta=r^{2}$ with $f(r)=1-{2GM}/{r}$. Here, $M$ denotes the mass of the BH. The effective gravitational potential perceived by the fluid is given by $\Phi^{\text{eff}}=\frac{1}{2}\ln f(r)$, and the radial momentum equation transforms into,
$v\gamma^{2}_{v}\frac{dv}{dr}+\frac{1}{h\tilde{\rho}}\frac{dp}{dr}+\frac{1}{2}\frac{f^{\prime}(r)}{f(r)}=0.$ (6)
The mass accretion rate is obtained from the continuity as,
$\frac{dM}{dt}=4\pi r^{2}\tilde{\rho}\gamma_{v}v\sqrt{f(r)}.$ (7)
Equations (6) and (7) are the two master equations that govern the spherical accretion dynamics for an asymptotically flat BH. However, to understand accretion onto PBHs, we reformulate these equations in the Friedmann–Lemaître–Robertson–Walker background, parametrized by the scale factor $a(t)$, with the metric given by $ds^{2}=-dt^{2}+a^{2}(t)(dr^{2}+r^{2}d\theta^{2}+r^{2}\sin^{2}\theta\,d\phi^{2})$. For this, we follow the conventional approach detailed in Ricotti:2007jk .
II.1 Accretion in the expanding background
The background expansion is incorporated by redefining the radial coordinate $r=a(t)x$. In this expanding coordinate, the radial velocity $v$ in terms of comoving coordinate $x$ and peculiar velocity $v_{p}$ becomes,
$v$ $=\frac{dr}{dt}=Hax+v_{p},$ (8)
where $H~{}(=\dot{a}/a)$ is the Hubble parameter and $v_{p}~{}(=a\dot{x})$. Incorporating Eq. (25) into the accretion equations, and integrating Eq. (6) we get,
$\ln\left(1-(Hax+v_{p})^{2}\right)+\frac{2\omega}{(1+\omega)}\ln\left(\frac{\rho_{\infty}}{\rho}\right)-\ln f(ax)=0.$ (9)
The integration constant is set by the condition that far away from the BH ($x\rightarrow\infty$), $v\rightarrow 0$, and $\rho\rightarrow\rho_{\infty}\propto a^{-3(1+\omega)}$, which corresponds to the background energy density of the Universe, characterized by the generic isothermal equation of state $p=\omega\rho$. The Hubble parameter satisfies the dynamical equation $H^{2}=\dot{a^{2}}/a^{2}=({1}/{3M_{p}^{2}})\rho$, which yields $H=H_{\rm in}\left(a/a_{\rm in}\right)^{-\frac{3}{2}(1+\omega)}$, with index “${\rm in}$” being some initial time which we identify later. Here, $M_{p}=1/\sqrt{8\pi G}\simeq 2.435\times 10^{18}\text{Gev}$, being the reduced Planck mass. Utilizing this, the relation between energy density and mass density of the fluid can be obtained by integrating the enthalpy generation Eq. (4) as,
$\tilde{\rho}^{1+\omega}=\rho_{\infty}^{\omega}\rho.$ (10)
To obtain this, we have assumed the mass and energy densities of the fluid at infinity are equal to $\rho_{\infty}$. Further, using Eq. (25) and Eq. (10), Eq. (7) can be written as,
$\frac{dM}{dt}=4\pi(ax)^{2}\rho_{\infty}^{\frac{\omega}{1+\omega}}\rho^{\frac{1}{1+\omega}}\frac{(Hax+v_{p})}{\sqrt{1-(Hax+v_{p})^{2}}}\sqrt{f(ax)}.$ (11)
The mass accretion rate $\dot{M}~{}(=dM/dt)$ is assumed to be a conserved quantity, and satisfies $d\dot{M}/dx=0$. Utilizing this and combining Eq. (9) and (11), one readily obtains the radial variation of the comoving fluid velocity as,
$\frac{dv_{p}}{dx}=\frac{\frac{2\omega}{(ax)}+\frac{\omega f^{\prime}(ax)}{2f(ax)}-\frac{f^{\prime}(ax)}{2f(ax)}-H\frac{v^{2}-\omega}{v(1-v^{2})}}{\frac{v^{2}-\omega}{v(1-v^{2})}\frac{1}{a}}=\frac{N}{D}.$ (12)
In Fig. 1, we have depicted the various representative solution topologies for the inflowing matter. However, the physical solutions are the ones where the inflowing matters pass through a critical point and transits smoothly from subsonic to supersonic branch. At the critical point, one needs to satisfy $N=D=0$ condition simultaneously. Solving this, we get the critical point location as $x_{c}=GM(1+3\omega)/(2a\omega)$ and matter velocity at $x_{c}$ as $v^{c}_{p}=\sqrt{\omega}-Hax_{c}$. Inserting these in Eq. (11), the accretion rate at the critical point becomes,
$\frac{dM}{dt}=\frac{\lambda_{c}}{16\pi M_{p}^{4}}M^{2}\rho_{\infty},$ (13)
where $\lambda_{c}=(1/4)\omega^{-3/2}(1+3\omega)^{(1+3\omega)/2\omega}$. Note that in the non-relativistic limit, the critical value reduces to $\lambda_{c}^{\rm NR}=(1/4)\text{exp}(3/2)\omega^{-3/2}$ Ali-Haimoud:2016mbv . Integrating Eq. (13), we obtain
$\frac{M}{M_{\mathrm{in}}}$ $=$ $\left[1-\frac{\lambda_{c}\gamma}{2(1+\omega)}\left(\frac{\frac{3}{2}H_{\rm in}(1+\omega)(t-t_{\rm in})}{\frac{3}{2}H_{\rm in}(1+\omega)(t-t_{\rm in})+1}\right)\right]^{-1},$ (14)
where $M_{\mathrm{in}}~{}(=4\pi\gamma M_{p}^{2}H_{\rm in}^{-1})$ is the initial mass of a PBH with $H_{\rm in}$ being the Hubble parameter at the formation time $t=t_{\rm in}$. Here, $\gamma~{}(=\omega^{3/2})$ is the collapse fraction to form PBH Carr & Hawking (1974). From Eq. (14), we indeed see that PBH gains mass due to accretion and quickly saturates into (see left panel of Fig. 2),
$M_{\rm acc}=M_{\mathrm{in}}\left(1-\frac{\lambda_{c}\gamma}{2(1+\omega)}\right)^{-1}.$ (15)
This is one of the key results of our analysis. For example, setting $\omega=1/3$ leads to a striking increase in the BH mass, reaching approximately $4.53\,M_{\mathrm{in}}$ [^1] Throughout our analysis, we assume non-rotating PBHs. However, if spinning PBHs are considered, the mass increase would be slightly higher—a possibility we leave for future investigation. The accretion time scale of attaining this mass can immediately be obtained from Eq. (14) at the time when accretion rate equals expansion rate, $i.e.$, when $\dot{M}/M=H$, which yields,
$t_{\rm acc}-t_{\rm in}=\tau_{\rm BH}^{\rm acc}\simeq\frac{3.24}{4\pi\gamma}\frac{M_{\rm in}}{M_{p}^{2}}.$ (16)
For example, if we consider a PBH of mass $10$ g accreting in the radiation background ($\omega=1/3$), accretion ceases at approximately $\tau_{\rm BH}^{\rm acc}\simeq 3\times 10^{6}M_{p}^{-1}$.
Figure 1: Plot of flow velocity ($v$) as a function of radial distance ($r$) for different accretion rate, with an equation of state $\omega=1/3$.
III Hawking evaporation
In the quantum mechanical framework, a BH can also emit particles via Hawking radiation. Due to this Hawking emission, BH loses its mass in the cosmological time scale. The BHs decay rate can be well described by Page:1976df ,
$\frac{dM}{dt}=-\epsilon(M)\frac{M^{4}_{p}}{M^{2}}.$ (17)
In the high energy limit of the emitted particles $E_{j}>>T_{\rm BH}$, called geometrical optical limit, the parameter $\epsilon(M)$ is approximately expressed as $\epsilon\simeq\frac{27}{4}\frac{g_{\ast}(T_{\rm BH})\pi}{480}$, with $g_{\ast}(T_{\rm BH})\sim 106.76$ being the number of relativistic degrees of freedom at the BH temperature $T_{\rm BH}=M_{p}^{2}/M$ Baldes:2020nuv . With this assumption, the PBH mass evolves as,
$M\simeq M_{\mathrm{in}}\left[1-\frac{(t-t_{\rm in})}{\tau_{\rm BH}^{\rm ev}}\right]^{\frac{1}{3}}.$ (18)
Similar to the accretion time scale, here also we have BH evaporation time scale set by $M=0$,
$t_{\rm ev}-t_{\rm in}=\tau_{\rm BH}^{\rm ev}\simeq\frac{M_{\rm in}^{3}}{3\epsilon M_{p}^{4}}.$ (19)
For example, assuming a PBH with a mass of 10 g formed in a radiation-dominated background ($\omega=1/3$), it completely evaporates at approximately $\tau_{\rm BH}^{\rm ev}\simeq 8.8\times 10^{17}M_{p}^{-1}$. This highlights the wide separation between the timescales of the two different physical processes of our interest, namely accretion and evaporation ($\tau_{\rm BH}^{\rm ev}\gg\tau_{\rm BH}^{\rm acc}$), as depicted in the left panel of Fig. 2.
The second natural quantity of our interest is the number of particles produced during its entire life time of a PBH. Following Haque:2023awl ; Haque:2024eyh ; Barman:2024slw , one can deduced the total number of emitted particles by a PBH of modified mass $M_{\rm acc}$, and temperature $T_{\rm BH}^{\rm acc}=T_{\rm BH}^{\rm in}\left(1-\frac{3\lambda_{c}\gamma}{8}\right)$ due to accretion as,
$N_{j}\simeq\begin{cases}10^{8}\left(\frac{M_{\rm acc}}{1\;\rm g}\right)^{2}~{}~{}~{}~{}~{}\text{for}~{}m_{j}\lesssim T_{\rm BH}^{\rm acc},\\ 10^{14}\left(\frac{10^{10}\text{Gev}}{m_{j}}\right)^{2}~{}\text{for}~{}m_{j}\gtrsim T_{\rm BH}^{\rm acc},\\ \end{cases}$ (20)
where ‘$j$’ stands for a particular emitted particle of mass $m_{j}$, and $T_{\rm BH}^{\rm in}~{}(=M_{p}^{2}/M_{\rm in})$ is the initial BH temperature. Note that Eq. (20) considers only scalar particles; for fermionic particles, a factor of $3/4$ should be included.
IV PBH mass evolution
In the preceding sections, we have discussed two important physical processes under which PBH evolves. Besides accretion being responsible for its increase in mass, a PBH also loses its mass by emitting Hawking radiation. Both the processes is described by the following PBH mass evolution equation as,
$\frac{dM}{dt}=\left(\frac{dM}{dt}\right)_{\text{accretion}}+\left(\frac{dM}{dt}\right)_{\text{evaporation}}.$ (21)
We note that while accretion takes place early in the PBH’s lifetime over a very short timescale $\tau_{\rm BH}^{\rm acc}$, evaporation occurs much later, near the end of its life. Due to this widely separated time scale of processes, we analytically solve Eq. (21) in two steps. As delineated earlier, in the first step, accretion enables the PBH to quickly gain mass, reaching $M_{\rm acc}=M_{\mathrm{in}}\left(1-\frac{\lambda_{c}\gamma}{2(1+\omega)}\right)^{-1}$. Accretion ceases at time $t_{\rm acc}-t_{\rm in}=\tau_{\rm BH}^{\rm acc}$, when the accretion rate equals the cosmic expansion rate, thereby setting the initial condition for the PBH’s subsequent evaporation phase. Upon solving the evaporation equation, we find that
$M\simeq M_{\rm in}\left[1-\frac{\lambda_{c}\gamma}{2(1+\omega)}\right]^{-1}\left[1-\frac{(t-t_{\rm acc})}{\tau_{\rm BH}}\right]^{\frac{1}{3}},$ (22)
where the PBH life time modifies into the following approximate form as,
$\tau_{\rm BH}\simeq\frac{M_{\mathrm{in}}^{3}}{3\epsilon M_{p}^{4}}\left[1-\frac{\lambda_{c}\gamma}{2(1+\omega)}\right]^{-3}.$ (23)
The evolution of PBH mass, accounting for both relativistic (blue) and non-relativistic (red) accretion in conjunction with Hawking evaporation, is illustrated in the left panel of Fig. 2. Results are presented for two initial masses, namely $10\,\mathrm{g}$ (dashed lines) and $10^{8}\,\mathrm{g}$ (solid lines), respectively.
V BBN Constraint
PBHs evaporating during the era of Big Bang Nucleosynthesis (BBN) can inject additional relativistic degrees of freedom, potentially altering the primordial element abundances predicted by BBN. To preserve the observed consistency of BBN results, PBHs must evaporate before the onset of BBN. This requirement immediately suggests that the PBH mass within the range of $10^{8}\,\text{g}\lesssim M_{\rm BH}\lesssim 10^{13}\,\text{g}$ Keith:2020jww should not form in the early universe. However, when PBH accretion is taken into account, these mass constraints are naturally altered.
In order to calculate the evaporation temperature ($T_{\rm ev}$), we set $\omega=1/3$ and rewrite the PBH lifetime (Eq. (23)) as,
$\tau_{\rm BH}\simeq\frac{1}{2H_{\rm acc}}\left(\frac{a_{\rm ev}}{a_{\rm acc}}\right)^{2}=\frac{M_{\mathrm{in}}^{3}}{3\epsilon M_{p}^{4}}\left[1-\frac{3\lambda_{c}\gamma}{8}\right]^{-3},$ (24)
where $H_{\rm acc}~{}(=4\pi\gamma M_{p}^{2}M_{\rm acc}^{-1})$ is the Hubble parameter at the end of accretion $t_{\rm acc}$. In this radiation dominated Universe with energy density $\rho_{r}=\frac{\pi^{2}}{30}g_{\ast}(T)T^{4}$, the scale factor and temperature are related as $a_{\rm ev}/a_{\rm acc}=T_{\rm acc}/T_{\rm ev}$, and hence, PBH evaporation temperature ($T_{\rm ev}$) is computed as,
$T_{\rm ev}\simeq\sqrt{\frac{3\epsilon}{2}}\left(\frac{\pi^{2}g_{\ast}}{30}\right)^{-1/4}\left(\frac{M_{p}}{M_{\mathrm{in}}}\right)^{3/2}\left(1-\frac{3\gamma\lambda_{c}}{8}\right)^{3/2}M_{p},$ (25)
where the relation $T_{\rm acc}=(\pi^{2}g_{\ast}/90)^{-1/4}(H_{\rm acc}M_{p})^{1/2}$ is used corresponding to radiation temperature at $H_{\rm acc}$. Combining Eqs. (24) and (25), the initial PBH mass can be expressed as a function of the evaporation temperature,
$\left(\frac{M_{\mathrm{in}}}{1\rm g}\right)\simeq 3.94\times 10^{8}\left(1-\frac{3\gamma\lambda_{c}}{8}\right)\left(\frac{T_{\rm ev}}{4\rm Mev}\right)^{-2/3}.$ (26)
In the above expression, we consider $g_{*}=106.75$ at the time of PBH formation. Equation (26) clearly demonstrates the impact of accretion on the BBN limit for the minimum PBH mass. For example, if we set $\lambda_{c}=0$ (without accretion), we recover the well known BBN bound on PBH initial mass $M_{\rm in}\simeq 3.94\times 10^{8}\,\mathrm{g}$. On contrary, relativistic accretion ($\lambda_{c}\simeq 10.4$) clearly shifts such bound to one order lower value, $M_{\rm in}\simeq 8.6\times 10^{7}\,\mathrm{g}$. Our analytic estimates match well with the numerical ones illustrated in the right panel of Fig. 2, and further depicts the dependence of the initial PBH mass on $T_{\rm ev}$ for different accretion scenarios.
Figure 2: Left panel: PBH mass ($M$) evolution as a function of scale factor ($a$) for initial masses of $10$ g (dashed) and $10^{8}$ g (solid), comparing relativistic (green) and non-relativistic (red) accretion along with Hawking evaporation. Right panel: Initial PBH mass ($M_{\rm in}$) as a function of evaporation temperature ($T_{\rm ev}$), shown for three cases namely, blue, red, and green curves representing without, with non-relativistic, and with relativistic accretion respectively.
VI Improved constraints on DM
PBH evaporation provides a compelling mechanism for explaining the observed DM abundance, $\Omega_{x,0}h^{2}\simeq 0.12$ Planck:2018vyg . Assuming that the present DM is produced completely from the evaporating PBHs, and subsequently no additional entropy has been produced, the DM abundance can be expressed as Cheek:2021odj ; Mambrini:2021cwd ,
$\Omega_{x,0}h^{2}\simeq 1.7\left(\frac{\gamma}{0.2}\right)^{\frac{1}{2}}\left(\frac{g_{\ast}}{106.75}\right)^{-\frac{1}{4}}\left(\frac{1\text{g}}{M_{\mathrm{in}}}\right)^{\frac{3}{2}}\left(\frac{m_{x}}{1\text{Gev}}\right)\beta N_{x},$ (27)
where $\beta=\rho_{\rm BH}^{\rm in}/\rho^{\rm in}$ is the PBH fraction defined as the ratio of the PBH energy density to the total background energy density at the time their formation. The total number of DM particles ($N_{x}$) produced from evaporation of PBH can then be computed using Eq. (20) for a given DM mass $m_{x}$. For our present purpose we consider scalar dark matter particle. For $m_{x}\lesssim T_{\rm BH}^{\rm acc}$, we have,
$\beta\leq 7\times 10^{-10}\left(1-\frac{3\lambda_{c}\gamma}{8}\right)^{2}\left(\frac{M_{\mathrm{in}}}{1\rm g}\right)^{-1/2}\left(\frac{1\rm Gev}{m_{x}}\right)$ (28)
and for $m_{x}>T_{\rm BH}^{\rm acc}$,
$\beta\leq 7\times 10^{-36}\left(\frac{M_{\mathrm{in}}}{1\rm g}\right)^{3/2}\left(\frac{1\rm Gev}{m_{x}}\right)^{-1}.$ (29)
Here, the equality represents the present day DM relic abundance shown in Fig. 3 using solid brown, cyan, green and magenta lines. The appreciable departure from the standard cases without accretion is indeed observed particularly for dark matter masses $m_{x}<T_{\rm BH}^{\rm acc}$ as shown in respective dotted lines. Blue lines correspond to critical $\beta_{c}\equiv T_{\rm ev}/T_{\rm acc}$, above which we have matter dominated universe, and $\beta$ becomes independent of $M_{\rm in}$ represented by Vertical lines. For comparison, we also recover the relation in the absence of accretion, shown by the dotted lines.
An important constraint on the DM parameter space comes from the warm dark matter (WDM) bound. PBHs can emit DM particles with high momenta, which may remain relativistic until structure formation, conflicting with observations. This bound is given by Haque:2023awl ; Barman:2024iht ,
$\frac{m_{x}}{\rm GeV}\geq 8.1\times 10^{7}\left(\frac{m_{\rm WDM}}{\rm keV}\right)^{\frac{4}{3}}\left(\frac{M_{\rm acc}}{M_{p}}\right)^{\frac{1}{2}}.$ (30)
In this analysis, we adopt the lower bound $m_{\rm WDM}>3.3\,\rm keV$ Viel:2013fqw ; Bode:2000gq . This constraint is shown in yellow in Fig. 3.
Figure 3: Variation of $\beta$ with $M_{\rm in}$ for various DM masses. Dotted lines represent scenarios with only Hawking evaporation, while the solid lines incorporate the effects of relativistic accretion as well. Shaded regions indicate constraints from inflation (light blue), BBN (green), improved BBN (light green), and Ly-$\alpha$ limits (yellow line).
VII Constrains on evaporating PBHs
PBHs with masses below approximately $10^{15}\,\rm g$ are expected to have completely evaporated via Hawking radiation by the present epoch in the absence of accretion effects Carr:2020gox ; Carr:2009jm . However, incorporating accretion into the PBH evolution notably extends their lifetimes, thereby modifying the mass bound for surviving PBHs today.
Assuming that PBHs evaporate at the present cosmic time $t_{0}\simeq 4.53\times 10^{17}\,\rm s$, their lifetime is approximately given by $\tau_{\rm BH}=t_{0}-t_{\rm acc}\simeq t_{0}$. Using this and the PBH lifetime relation in Eq. (23), the modified initial mass for PBHs evaporating today is estimated as,
$M^{0}_{\rm in}\simeq\left(3\epsilon M_{p}^{4}t_{0}\right)^{\frac{1}{3}}\left(1-\frac{3\gamma\lambda_{c}}{8}\right)\simeq 1.2\times 10^{15}\rm g\left(1-\frac{3\gamma\lambda_{c}}{8}\right).$ (31)
This immediately modifies the initial PBH mass limit of approximately $\simeq 2.72\times 10^{14}\,\rm g$, below which all PBHs would have fully evaporated by now. In the absence of relativistic accretion ($\lambda_{c}=0$), this limit is about an order of magnitude higher, at $1.24\times 10^{15}\,\rm g$. The inclusion of accretion thus introduces a new threshold, naturally shifting the parameter space of present-day PBH abundance, parametrized by $f_{\rm PBH}(M)=\Omega_{\rm PBH}/\Omega_{\rm DM}$, as a function of their initial mass $M_{\rm in}$. Here, $\Omega_{\rm PBH}$ and $\Omega_{\rm DM}$ denote the present density parameters of PBHs and total DM, respectively. The constraints on $f_{\rm PBH}(M)$ arise from several observational channels Oncins:2022ydg ; Carr:2016drx ; Green:2020jor ; Carr:2020xqk , including gravitational lensing Niikura:2019kqi ; Niikura:2017zjd , accretion-induced emissions Manshanden:2018tze ; Ali-Haimoud:2016mbv , extragalactic gamma-ray backgrounds Carr:2016hva , and gravitational wave observations Nitz:2021vqh ; Kavanagh:2018ggo . Since accretion alters the mass evolution of PBHs, it induces a noticeable change in the viable parameter space, as illustrated in Fig. 4 using solid curves. This evidently leads to the modified mass range within $10^{16}\,{\rm g}$ to $10^{21}\,{\rm g}$ are still allowed to serve as the sole DM candidate.
Figure 4: Constraints on the PBH dark matter fraction $f_{\rm PBH}$ vs. initial mass $M_{\rm in}$ are shown. The dashed curve includes only Hawking evaporation; the solid curve also accounts for relativistic accretion. Shaded regions denote bounds from evaporation (red), microlensing (blue), and GWs (purple). Black lines mark the critical mass with (solid) and without (dashed) accretion.
VIII Gravitational Waves from PBH Evaporation
PBHs evaporate via Hawking radiation, emitting gravitons that contribute to a stochastic gravitational wave (GW) background. This GW signal offers a unique window into PBH physics and the early universe. Accurate modeling requires going beyond idealized blackbody assumptions using tools like BlackHawk Arbey:2019mbc , which accounts for greybody factors, spin effects, and realistic emission spectra.
The instantaneous GW energy density is given by  Anantua:2008am ; Ireland:2023avg ,
$\frac{d\rho_{\text{GW}}}{dt\,d\omega}=n_{\text{BH}}(t)\,\frac{\omega}{2\pi}\,Q_{\text{GW}}(t,\omega,M_{\rm acc}),$ (32)
where $Q_{\text{GW}}$ is the graviton flux computed numerically. The GW energy density at evaporation is obtained by integrating over the PBH lifetime; redshifting this result gives the present day spectrum as,
$\Omega_{\rm GW}(f)=\frac{\rho_{\rm BH}^{\rm in}}{M_{\rm in}\,\rho_{\rm c}^{0}}\frac{\omega_{0}^{2}}{2\pi}\frac{a_{\rm in}^{3}}{a_{0}^{2}}\int_{t_{\rm acc}}^{t_{\rm ev}}\frac{dt}{a(t)}\,Q_{\text{GW}}\left(t,\frac{\omega_{0}a_{0}}{a(t)}\right),$ (33)
where $\rho_{\rm c}^{0}$ and $a_{\rm 0}$ denote the critical energy density and scale factor measured today, respectively.
Figure 5: $\Omega_{\rm GW}$ is shown versus frequency. The solid line includes both Hawking evaporation and relativistic accretion; the dashed line shows only Hawking emission. Red curves indicate $\Delta N_{\rm eff}=0.17$ (Planck + BAO)  Planck:2018vyg and $\Delta N_{\rm eff}=0.06$ (CMB-S4)  Abazajian:2019eic constraints.
Relativistic accretion extends PBH lifetimes by modifying the mass loss rate, which in turn shifts the peak frequency and amplitude of the resulting GW spectrum (see Fig.  5). Detecting these high-frequency GWs is challenging but potentially feasible with next-generation detectors, including laser interferometers (1–10 kHz)  Ackley:2020atn , levitated sensors (10–100 kHz)  Aggarwal:2020olq , magnetic conversion (10 GHz)  Li:2009zzy ; Ito:2019wcb , and inverse Gertsenshtein effect devices ($10^{14}–10^{18}$ Hz)  Ejlli:2019bqj . Axion experiments such as JURA Beacham:2019nyx , OSQAR II OSQAR:2015qdv , and CAST  CAST:2017uph may be repurposed for graviton detection, providing novel probes of PBH accretion and early-universe dynamics.
IX Conclusions
PBHs have become an integral part of the theoretical model building of the early Universe and dark matter. In this Letter, we revisit the cosmological evolution of PBHs by incorporating two key physical processes, namely Hawking evaporation and accretion. In contrast to previous studies that modeled accretion using a pseudo-Newtonian approximation, typically yielding negligible mass growth, we formulate the accretion dynamics within a fully general relativistic framework. We find that relativistic corrections significantly enhance the accretion rate, leading to rapid PBH mass growth on timescales much shorter than those of Hawking evaporation. This enhancement imposes a new lower bound on the PBH mass, $M_{\rm in}\leq 6.9\times 10^{7}$ g, above which excessive energy injection from evaporation can disrupt BBN.
Furthermore, we analyze the consequences of this enhanced accretion on DM production via Hawking radiation and find a noticeable shift in the allowed parameter space spanned by PBH abundance and DM mass, modifying constraints relative to previous results Cheek:2021odj ; Gondolo:2020uqv . We also determine that the minimum PBH mass capable of surviving to the present day is reduced to $\sim 2\times 10^{14}$ g, which is an order of magnitude below existing estimates Carr:2009jm ; Carr:2020xqk . PBHs with masses above this threshold could themselves constitute viable DM candidates. Notably, a broad mass window between $10^{16}\,{\rm g}$ and $10^{21}\,{\rm g}$ remains observationally unconstrained, allowing PBHs within this range to account for the entirety of DM. Finally, we compute modifications to the high-frequency GW spectrum arising from accretion-driven PBH mass growth, offering a potential observational signature for future detectors.
X ACKNOWLEDGEMENTS
MRH acknowledges ISI Kolkata for providing financial support through the Research Associateship. JK would like to thank Gargi Sen and Samik Mitra for their help and useful discussions. DM gratefully acknowledges support from the Science and Engineering Research Board (SERB), Department of Science and Technology (DST), Government of India (GoI), through the Core Research Grant CRG/2020/003664.